jeremaiah is now 10 yrs older than his sister justin. in five years jeremiah's age will be twice justine's age. how old are they now?
Answer by checkley71(8403)

JUSTINE=X NOW
JEREMIAH=X+10 NOW
IN 5 YEARS JUSTINE WILL BE X+5
IN 5 YEARS JEREMIAH WILL BE X+10+5=X+15
2(X+5)=(X+15)
2X+10=X+15
2X-X=15-10
X=5 AGE OF JUSTINE NOW.
5+10=15 AGE OF JEREMIAH NOW.
PROOF
2(5+5)=5+15
2*10=20
20=20